# When Gaussian Meets Surfel: Ultra-fast High-fidelity Radiance Field Rendering

### v1

@@ Appendix B Number of Primitives
In Table 6, we report the number of primitives in an outdoor scene (Garden) and an indoor scene (Room) of different methods. Our surfel pruning strategy efficiently prunes a considerable amount surfels, and only a small number of surfels are retained to produce coarse geometry and appearance. The total number of primitives in 3D-GES is close to that of AdrGS (Wang et al., 2024), but 3D-GES achieves faster frame rates and higher rendering quality, which further demonstrates the efficiency of GESs.
@@ Appendix D More Comparisons
In Fig. 20, we show more comparisons on geometry reconstruction quality between 2D-GES and 2DGS (Huang et al., 2024). With the surfels serving as the primary geometry and ensuring multi-view geometry consistency, our method can reconstruct more robust geometry. Meanwhile, the Gaussians in 2D-GES not only smooth the geometry but also introduce additional geometry details.

### v2

@@ Appendix B Number of Primitives
In Table 6, we report the number of primitives in an outdoor scene (Garden) and an indoor scene (Room) of different methods. Our surfel pruning strategy efficiently prunes a considerable amount surfels, and only a small number of surfels are retained to produce coarse geometry and appearance. The total number of primitives in 3D-GES is close to that of AdrGS (adr2024), but 3D-GES achieves faster frame rates and higher rendering quality, which further demonstrates the efficiency of GESs.
@@ Appendix D More Comparisons
In Fig. 20, we show more comparisons on geometry reconstruction quality between 2D-GES and 2DGS (2dgs2024). With the surfels serving as the primary geometry and ensuring multi-view geometry consistency, our method can reconstruct more robust geometry. Meanwhile, the Gaussians in 2D-GES not only smooth the geometry but also introduce additional geometry details.